### v3

# KeySpace: Public Key Infrastructure Considerations in Interplanetary Networks

Abstract
As satellite networks expand to encompass megaconstellations and interplanetary communication, the need for effective Public Key Infrastructure (PKI) becomes increasingly pressing. This paper addresses the challenge of implementing PKI in these complex networks, identifying the essential goals and requirements.
We develop a standardized framework for comparing PKI systems across various network topologies, enabling the evaluation of their performance and security. Our results demonstrate that terrestrial PKI techniques can be adapted for use in highly distributed interplanetary networks, achieving efficient low-latency connection establishment and minimizing the impact of attacks through effective revocation mechanisms. This result has significant implications for the design of future satellite networks, as it enables the reuse of existing PKI solutions to provide increased compatibility with terrestrial networks.
We evaluate this by building the Deep Space Network Simulator (DSNS), a novel tool for efficiently simulating large space networks. Using DSNS, we conduct comprehensive simulations of connection establishment and key revocation under a range of network topologies and PKI configurations. Furthermore, we propose and evaluate two new configuration options: OCSP Hybrid, and the use of relay nodes as a firewall. Together these minimize the extent of the network an attacker can reach with a compromised key, and reduce the attacker’s load on interplanetary relay links.
1 Motivation
Figure 1: Example of an interplanetary network of satellites and the communication between them. There is a federated constellation of satellites around Earth (left), and relays connect the network to the Moon/Mars (right). These links are high-latency and, due to satellite and planet movement, are not always available.
Due to the decreasing cost of launching and operating satellites, the number of satellites in orbit has grown dramatically in recent years, and is only set to increase with time. In particular, the availability of commodity components and rideshared launches have made it especially easy to launch CubeSats – these satellites have a small form factor and well-defined dimensions so that many satellites can fit on a single launch vehicle. As this trend continues, there will be a huge number of satellites around Earth in a wide range of orbits, as well as around other bodies in our solar system like the Moon and Mars, or in deeper space [israelLunaNet2020, nasascienceMars]. This raises some interesting questions surrounding communication: how do we ensure communication with these satellites is possible (to relay data for telemetry, control, and normal operation), and how do we ensure this communication is secured via authentication or authenticated encryption?
Two main architectures for large-scale satellite communication are emerging: relay networks and federation. The first of these enables communication between nodes in a network by relaying traffic over a small number of high-bandwidth links. This has already seen significant use by organizations like NASA and ESA using their respective relay satellite systems [monaghanTracking2015, esaEDRS2022], with upcoming proposals like NASA’s LunaNet and ESA’s Moonlight set to provide additional relays around the moon for lunar communication [israelLunaNet2020, nasaLunaNet2022, europeanspaceagencyMoonlight2024]. Federated Satellite Systems (FSSs) provide additional communication redundancy by enabling satellites to communicate with one another over Inter-Satellite Links (ISLs) beyond the constraints of traditional operator boundaries. There has been some recent interest in this concept, with works showing that such a network is feasible and useful, and beginning to consider protocols for negotiating federations on-the-fly to share bandwidth or payloads [aranitiOpportunistic2021, lluchSimulating2015, ruiz-de-azuaProofofconcept2019, ruiz-de-azuanovel2020, ruiz-de-azuaFederated2021]. Supporting both of these is the increased deployment of terrestrial communication systems, with projects like Amazon’s AWS Ground Station providing centralized control of satellites with broad coverage, using a network of ground-level antennas [amazonwebservicesAWS2023].
This results in a network topology that has a number of sporadic, high-latency links (see Figure 1). The network naturally partitions itself based on these long-distance links, with good connections within each partition. This type of network is considered to be a Delay-Tolerant Network (DTN), an architecture which has been explored extensively in related works. In DTNs, key management is considered to be an unsolved problem [templinDelay2014] and is an active area of research, with the majority of research focusing on the unpredictable nature of the network, making use of techniques like gossip protocols and web of trust to provide probabilistic assurances of success despite the network’s unpredictable movement and connectivity [khareWeaving1997, garfinkelPGP1995, eschenauerKeyManagement2002, menesidouCryptographic2017, deandradeFully2016].
However, although connectivity in satellite networks is intermittent, the entire network topology is predictable, including the movement of satellites and availability of links. This differentiates satellite networks from other types of DTNs; as a result the underlying assumptions of key management in DTNs break down. Data does not need to be routed probabilistically and can instead be optimally routed to its destination, even if the network topology changes in-transit. In contrast to current consensus, we thus hypothesize that we can apply existing terrestrial Public Key Infrastructure (PKI) concepts.
PKI is well-suited to future satellite networks, due to its built-in hierarchical nature (allowing operators to delegate authority while retaining centralized control) and effective mechanisms for verifying identity and validity. However, mechanisms for updating and revoking keys are centralized and rely upon distribution of revocation lists or low-latency communication with certificate authorities; for this reason they have been largely dismissed for use in DTNs, in which low-latency connections cannot always be guaranteed, and broadcasts are probabilistic [templinDelay2014, bhuttaSecurity2009, sethPractical2005].
Contributions. In this paper, we outline the goals and requirements for deploying and operating PKI in large satellite networks, with a particular focus on interplanetary networks with long-distance, intermittent relays. We construct a standardized set of experiments for comparing different key management systems to one another for a given network topology, looking in particular at the performance of connection establishment procedures, and the speed at which key revocations and updates can occur so that the damage caused by a compromised key can be minimized. We evaluate these through the use of simulations, building the Deep Space Network Simulator (DSNS): an extensible system for the realistic simulation of large networks of satellites. Through this analysis we demonstrate, for the first time, that commonly used terrestrial PKI protocols can be employed in satellite networks. We also establish configurations which optimize performance and security, and propose novel variations on these protocols which provide more effective revocations to protect remote segments more quickly. We do not compare to novel protocols designed for DTNs; rather, we demonstrate that existing protocols can be effectively used in satellite networks, thereby maintaining compatibility with terrestrial networks.
2 Background
In this section we outline key concepts used in this paper, including an overview of key management in DTNs and space systems.
Delay Tolerant Networks (DTNs). DTNs encompass a wide range of networks including mobile ad-hoc networks, wireless sensor networks, and networks in space [torgersonDelayTolerant2007]. These networks are characterized by a lack of persistent connectivity and unpredictable network topology. This requires changes to the network stack, particularly in routing: end-to-end routes cannot always be precomputed, or even established in the first place, due to the structure of the network. Instead, standards like the Bundle Protocol (BP) are used, enabling store-and-forward delivery [scottBundle2007], and proposed routing techniques often rely upon probabilistic methods [grasicEvolution2011, leguayDTN2005].
However, unlike other types of DTN the movement of satellites is fixed, so connectivity can be predicted ahead of time with near perfect precision. As a result, optimal routes can be computed ahead of time using techniques like Contact Graph Routing (CGR) [aranitiContact2015]; however, store-and-forward delivery is still required due to the sporadic nature of links. Sometimes referred to as the Interplanetary Internet (IPN), there exists a good body of work addressing networking under these particular constraints [akyildizState2004, elalaouiMARS2020, alhilalSky2019], including recommendations for architecture and governance from the Internet Society [ipnsigSolar2023], and protocol and architecture recommendations from the Consultative Committee for Space Data Systems (CCSDS) [ccsdsReference2008, ccsdsOverview2014].
Public-Key Infrastructure (PKI). A PKI attaches identities to public-private keypairs to secure communication in larger networks. It does this through certificates provided by Certificate Authorities (CAs), which verify a keypair’s identity. A network participant can check a keypair’s legitimacy by verifying its certificate, signed by a trusted CA which is part of a signature chain established by the root CA. This system is near-universal in the terrestrial internet, using the X.509 standard [boeyenInternet2008].
To verify a certificate’s trustworthiness, the recipient must check that it has not been revoked. This can be achieved using Certificate Revocation Lists (CRLs), listing all revoked certificates, or using the Online Certificate Status Protocol (OCSP), which allows devices to query the CA about the status of a certificate [santesson5092013]. A variant called OCSP Stapling shifts the burden to the sender, who queries the CA and attaches the response to the message, reducing the risk of denial-of-service attacks and allowing for faster communication over long-distance links [eastlake3rdTransport2011].
However, in DTNs both CRLs and OCSP are largely dismissed as infeasible: CRLs are too bandwidth-intensive and OCSP requires constant low-latency access to CAs, which cannot always be guaranteed. Instead, DTNs often make use of gossip protocols for key distribution – nodes exchange information with their peers in the hope that this information spreads as far as possible [eschenauerKeyManagement2002]. In this paper we demonstrate that these assumptions do not hold for all DTNs, and that the predictable links and network topology of interplanetary networks enables many terrestrial PKI concepts to be effective. Furthermore, once the above arguments have been addressed it can be argued that PKI is ideal for space applications, scaling particularly well to huge numbers of nodes, thus future-proofing space networks as they continue to grow even larger.
Current Systems. Many current satellite systems opt to use pre-shared encryption keys. This is significantly simpler to deploy in small systems, as keys can be loaded onto the satellite before launch, but does not scale well to systems with more satellites or shared operation. It also presents additional risks in cases where keys are leaked, broken, or reverse engineered (see the GK-2A and COMS-1 satellites [xrit-rx, lrit-key-dec]). Some systems support over-the-air rekeying in this case, using tightly controlled keys to authenticate the operation of replacing session keys [andersonAuthentication2023, ccsdsSpace2020].
The majority of current satellite constellations are privately owned (e.g., Starlink), so less is known about their internal operation. The Iridium constellation uses pre-shared keys stored in the SIM card, with no mechanism for updating keys short of rewriting or replacing the SIM [veenemanIridium2021]. The key management of the satellites themselves is not known.
3 Related Work
Due to its wide range of applications, the issue of security in DTNs and satellite networks has been well-explored, both in scientific literature and in internet drafts and RFCs.
3.1 Internet Drafts
The most recent proposed standards for the Bundle Protocol include the BPSec security extensions [burleighBundle2022, birraneBundle2022], which add options for integrity and confidentiality to secure BP communication. Notably, this standard does not recommend a specific method for key management, assuming it to be handled separately [birraneBundle2022]. In contrast, in this paper we focus on the key management mechanism, assuming the underlying cryptosystem to be provided through BPSec or a similar protocol.
Other drafts assess the problem of key management in DTNs and space networks. The authors of [farrellDelayTolerant2009] recognize key management to be an important open issue in DTNs – this is expanded on in [templinDelay2014], which argues that the long delays and expected disruption in DTNs makes traditional PKI methods unsuitable due to their requirement of short-turnaround communication with CAs.
In [farrellDTN2007], a basic set of high-level requirements is given for key management systems in DTNs; we draw particular attention to their note that systems must account for nodes with highly restricted connectivity, computation, or storage capacity. The authors of [templinDTN2016] also set out requirements for a PKI in DTNs, noting that the system must not rely on time-constrained queries to a central authority or have a single point of failure, and that revocations must be delay tolerant.
In [ccsdsSpace2011], the Consultative Committee for Space Data Systems (CCSDS) give a high-level overview of many potential key management concepts in space DTNs. There is also an ongoing CCSDS project to design a cloud-based CA for satellite systems, to be used as a common route of trust particularly in shared systems [ccsdsProject2021]. Results from our simulations will help to inform this and other projects – this is discussed further in Section 8.
In [viswanathanArchitecture2016], the authors lay out five potential design patterns for PKI in DTNs, which we use as a basis for some of our simulations. The “request-response” pattern describes OCSP and OCSP Stapling, “publish with proxy subscribe” matches behavior seen in OCSP with Validators, and “blacklist broadcast” resembles CRLs.
3.2 Academic Works
Reference [menesidouCryptographic2017] provides a good overview of existing works looking at key management in DTNs. Some of these focus on space networks or are network-agnostic, but the majority look at less relevant applications of DTNs such as rural networks.
A number of works focus on security initialization or authenticated key exchange – that is, initiating a secure channel where one did not previously exist by exchanging or establishing keys [menesidouAuthenticated2012]. We consider these to be out of scope of this paper, and focus on communication over (and updates within) an already established network.
The authors of [bhuttaSecurity2009] recognize that terrestrial protocols do not perform well in DTNs and smaller satellite networks, due to their reliance on constant connectivity and low latency communication. They also acknowledge that any key management system will need to account for the types of nodes and heterogeneity in the network. In this paper we provide a simulator that can be used to assess the differences between satellite networks, and set out guidelines that apply to satellite systems in general.
In [vonmaurichData2018] the authors focus on real-world assessments of various cryptosystems on embedded hardware. Additionally, they discuss the requirements of security protocols in federated satellite systems. They propose a high-level PKI-based protocol to guarantee confidentiality, integrity, and authenticity, and rely on a centralized authority to verify and distribute keys.
Some works such as [johnsonProviding2013] adapt PKI for use in satellite networks – their system is very similar to OCSP stapling systems discussed above, with nodes requesting an authentication pass from a CA ahead of communication. On the other hand, [sethPractical2005] argues that PKI is not suitable for use in DTNs, instead suggesting the use of Hierarchical Identity Based Cryptography (HIBC), in which keys are derived from public information about nodes, and a hierarchy of authorities manages keys. This is argued to be better than traditional PKI, as the public parameters do not need to be verified as frequently [asokanSecuring2007], at the cost of requiring additional secrets to be distributed out of channel. We argue that this is not necessary for satellite networks, as each network segment is sufficiently connected that it is always possible to query a trusted authority, provided there are sufficiently many distributed throughout the network. We demonstrate this using simulations in Section 7. HIBC also raises concerns about large-scale deployment; distributing secrets and maintaining identities out-of-channel becomes significantly more difficult as the number of users grows.
Reference [deandradeFully2016] proposes a decentralized key management system by passing around chains of certificates, with nodes added to the chain when another node trusts it. This has the benefit of being fully decentralized, but has no mechanisms for key revocation or centralized/federated management.
Finally, some works extend existing PKI concepts to make them more suitable for use in DTNs. The authors of [cooperMore2000] and [bhuttaPublic2017] build upon traditional CRLs, reducing the storage and complexity requirements by sending “delta CRLs” containing incremental updates and by using hash tables. In [koisserTruSat2022], the authors also extend PKI concepts, building a consensus protocol to remove the need for a single root CA, and testing a number of revocation methods including OCSP Stapling and Bloom filters. The authors of [luoDICTATE2005] recognize that a single centralized CA is not enough for PKI in DTNs, and propose the use of distributed CAs to ensure reachability. In [fangAdaptive2013] this concept is extended to “adaptive CAs”, in which any node can become a CA, and decisions are made dynamically.
3.3 Simulators
Some of these works are evaluated using simulators such as ns-3 [rileyns32010], the ONE Simulator [keranenONE2009], and OMNeT++ [vargaOMNeT2010]. However, these simulators each come with their downsides – ns-3 and OMNeT++ do not scale well to large numbers of nodes with lots of connections, and simulate the entire network stack, which is not necessary in many cases. The ONE Simulator is designed with DTNs in mind and is significantly more lightweight, but it is no longer maintained and does not natively support simulating satellite mobility, or the protocols we intend to evaluate. There also exists an extension to ns-3 for simulating satellite networks [kassingExploring2020], but due to the limitations of ns-3 it does not support dynamic links based on proximity or other factors. It is therefore limited to a subset of the satellite network configurations we evaluate in this paper.
In evaluations of DTN protocols, it is standard to simulate nodes moving randomly around a space, connecting to nearby nodes (for example, [djamaludinRevocation2016]) – whilst this is sufficient for some terrestrial DTNs, it does not capture the interesting mobility properties of a satellite network. In this paper we ensure our simulation system is designed such that constellation topology and connectivity can be simulated as realistically as possible. This allows us to perform a convincing real-world evaluation of key management in satellite systems, mirroring real-world deployments and proposals for future systems.
4 Threat Model
In this paper we consider an attacker that has gained access to the private key of a legitimate user on the network, and has at least one node in the network under their control. Using this key they can craft messages as if they were the legitimate user. There are a number of ways this could occur: compromising a satellite (e.g., by compromising telecommand authentication, taking control of a satellite via a compromised ground segment, or via a malicious hosted payload), intercepting improperly encrypted messages, compromising the supply chain, etc.
At a high level, the goal of the attacker is to cause damage within the network by using the compromised key to communicate with as many nodes as possible, for as long as possible. This will look different depending on the nature of the network and the compromised device: for example, producing forged measurement data, impersonating users in communication, or forging telecommand packets to damage the device itself.
When the attacker is discovered, the compromised key must be revoked. The long distances between nodes in an interplanetary network and inherent speed-of-light limit means the goal is not to prevent the attacker entirely, but instead to minimize the extent of the damage from the moment the attack is identified. This is achieved by propagating the revocation as efficiently as possible.
5 Framework Design
In this section we establish the goals and requirements for key management in satellite networks, constructing from them a set of standardized experiments which can be used to compare protocols against one another. Rather than being limited to one network topology or configuration, these are instead designed to be high-level tests that can be used on any network to compare different protocols side by side. This allows network operators to test protocols in the context of their own network configuration, enabling well-informed decision making about which configuration to choose, balancing performance against security. In Section 6 we design experiments using standard terrestrial PKI protocols in interplanetary satellite networks, and evaluate improvements tailored to the unique topology of these networks.
We state the following goals for PKI in satellite networks:
  • Low latency: The latency of messages should be as close as possible to the shortest path between the two nodes. For many PKI systems this will not be a factor, but configurations which require messages to be sent via an authority node should be configured to minimize the additional distance traversed by the message.
  • Low establishment overhead: Establishment overhead $T_{E}$ is computed to be the time difference between sending a message from one node to another, compared to opening a connection between the two nodes (i.e., a full handshake). The long distances traversed by messages between planetary segments, combined with the short duration some relay links may be available, makes minimizing this figure crucial.
  • Fast revocation coverage: In large satellite networks with many users, it is inevitable that keys will need to be updated and revoked. We consider the revocation coverage time $T_{R}(s)$ to be the time at which all of the nodes in segment $s$ are protected due to the revocation information having propagated sufficiently far (e.g., to the relevant CA). There will be a necessary lower bound to this figure, but protocols and network configuration should be tweaked to optimize this figure.
  • Low attack penetration: In cases where authentication data has been cached, we also evaluate the penetration of an attack $P(s)$ to be the proportion of nodes in segment $s$ which accept an attacker’s message sent at the same time as a revocation was issued. This must be minimized where relevant.
We also propose the following secondary goals, which are not required but may be desirable:
  • Distributed: In order to achieve good performance across long distances, authority should be distributed throughout the network to enable querying without excessive latency. Later on we discuss the trade-off between centralized control and performance.
  • Intercompatible: We argue that in order for interplanetary connectivity to be effective, protocols will need to work alongside existing internet infrastructure. Protocols may be designed to work seamlessly with existing protocols, or utilize a translation layer. Protocols should also be able to be made compatible with BPSec [birraneBundle2022], which provides mechanisms to secure communication but leaves key management as an open question.
Our focus is on adapting PKI mechanisms to the environment of interplanetary networks, rather than on the underlying cryptosystem itself. While considerations such as storage, computational requirements, and bandwidth usage are crucial in this context, they are inherently tied to the choice of cryptosystem and are therefore outside the scope of this work. By concentrating on the feasibility of applying terrestrial PKI mechanisms in space, we establish a foundation for future research, which can then focus on optimizing the cryptosystem and its components to meet the unique demands of interplanetary networks.
5.1 Scenarios
We construct two scenarios to evaluate the above goals: connection establishment, and key revocation.
5.1.1 Connection Establishment
In the first scenario one node wants to communicate with another, but must first prove its identity. To do so it sends a signed message with a certificate attached. The recipient validates the certificate via validity information either stapled to the message or requested from a CA.
This adds a time delay before communication can occur, which can be quite significant in satellite networks with long-distance links. This is a particular concern when communication occurs over a sporadic link between network segments; any time spent establishing the sender’s identity is time in which commmunication cannot occur, taking away from the total duration of the link’s uptime.
[…]
5.1.2 Key Revocation
In the second scenario we consider a key in use by a malicious actor. The key is revoked, and we aim to minimize the penetration of the attack beyond this point, and ensure the revocation propagates across the network as quickly as possible.
Coverage time $T_{R}(s)$ (for segment $s$) is assessed relative to the revocation time $t_{0}$ by finding the time at which each segment’s CA has received the revocation message, adjusting to account for query time, and taking the greatest such time within the segment. As a result, for certain segments the coverage time may be negative – if a revocation reaches the CA quickly, but it takes a long time for the attacker’s message to reach that same CA (for instance, if there is a long-distance link between them), then nodes under that CA may be protected even from messages initialized before the revocation was started. Also note that in some cases such as OCSP Stapling the coverage will depend on the attacker’s location/CA; in these cases, results are gathered for every possible attacker location and aggregated by segment. We also assess attack penetration $P(s)$ for cases where certificate validity information is cached, with a fixed attacker location chosen to be in an opposite orbital shell from the CA in segment $s$.
6 Experiment Design
We aim to evaluate the feasibility of terrestrial PKI techniques on interplanetary networks using the framework established above, by testing a range of these techniques on a number of different network topologies. In this section we outline the key management systems and network topologies tested, and establish any additional experimental variables and constraints. Finally, we present a new network simulator to execute these experiments.
6.1 Key Management Systems
In our experiments we evaluate a number of PKI systems based on the existing and proposed mechanisms discussed in Section 3. In each case, nodes have a public/private keypair which is used to sign and encrypt messages. These keypairs are accompanied by a certificate – that is, a chain of signatures tracing back to a Certificate Authority (CA) vouching for its legitimacy. We specifically consider the following systems to verify a certificate’s validity, ensuring it has not been revoked:
  • Certificate Revocation Lists (CRLs): CAs each maintain a list of revoked certificates, which are sent to nodes on request. Any certificates on this list are rejected, and others are accepted provided there is a trusted CA in the certificate chain. CRLs are cached, with a new CRL requested on the next message received after expiry. To reduce size, CRLs can be built using hash tables [bhuttaPublic2017] or using delta CRLs [cooperMore2000].
  • CRL Broadcast: Similar to the above, but CAs broadcast CRLs at regular intervals, instead of on request.
  • Online Certificate Status Protocol (OCSP): On receipt of a signed message, nodes request the certificate’s validity information from their CA. The response is cached to reduce the overhead on subsequent messages.
  • OCSP Stapling: Similar to the above, but the node sending the message requests its CA for a signed message with validity information. This is sent alongside the original message, reducing the burden on the recipient.
  • OCSP with Validators: Proposed in [viswanathanArchitecture2016], messages are instead sent via a CA, which staples the certificate validity to the message in-transit. This removes the need for a full handshake in certificate validation, and can significantly reduce latency if the CA is en route to the message destination. This shares some similarities with the proposed “Revocation in the Middle” technique, using middleboxes to provide revocation information [szalachowskiRITM2016].
  • OCSP Hybrid: We propose a new system, combining and extending the above OCSP concepts. We note that in-transit OCSP validation increases congestion in a lot of cases by sending all traffic via the CA, but makes a lot of sense when traffic is traveling via a small set of relays anyway (e.g., traffic between Earth and Mars). We therefore propose a hybrid system in which standard OCSP Stapling is used for traffic within a segment, but any cross-segment traffic uses in-transit validation, with CAs located at the relays between network segments.
Sequence diagrams for each protocol are in Appendix B.
We evaluate each of these under a “centralized” model, with a single terrestrial CA, and a “distributed” model. In this case, there is one CA per network segment (i.e. layer of satellites or ground systems), which are designated as sub-CAs of a single terrestrial root CA. Authority is distributed such that any CA can provide validity information for any other certificate. This has the potential to significantly reduce the latency overhead of validating cross-segment messages, although the threat of attackers using revoked certificates from other segments is also increased. This motivates the revocation case, in which we demonstrate protection from these attacks with proper configuration.
In the revocation scenario, we also consider the case in which nodes have already cached validity information for the compromised certificate – this enables us to see the difference in effect between the situation where attacker messages are accepted by default, and one in which up-to-date revocation information must be retrieved. It also enables us to consider an alternative PKI configuration; in OCSP we have the option of subscribing nodes to updates, with CAs remembering which nodes have been sent certificate validity information and passing on certificate updates to the relevant nodes. This allows revocations to take place more quickly without requiring broadcasts across the whole network. [^1] This option can be used for each OCSP configuration except OCSP Stapling – the CA does not know where the certificate status will be sent, so it cannot subscribe the recipient to updates.
Finally, we propose and evaluate a new technique: using interplanetary relay nodes as a firewall. We note that in interplanetary networks all cross-segment traffic must travel via relays, providing an excellent opportunity to remove messages from revoked keys before they reach their destination. This can significantly reduce load on the relay links by dropping attackers’ messages much sooner, and may also bring earlier revocation coverage, reducing the number of nodes reachable by an attacker with a revoked key.
6.2 Network Topologies
TABLE I: The number of nodes in each of the network topologies used in the experiments in this paper. Relays and multiple segments are separated by “+”.
Earth Moon Mars Relay
Network Ground Space Ground Space Ground Space
Earth/Moon/Mars 256+3 66 12 8 12 66 2
Earth/Mars 256+3 66 12 66 1
Earth/Moon 256+3 66 12 8 1
Iridium 256 66
Starlink 256 1584
CubeSat 256 121
LEO/LEO 256 66+1584
LEO/MEO 256 66+13
LEO/GEO 256 66+3
LEO/MEO/GEO 256 66+13+3
We want to ensure our experiments remain as true to real-world networks as possible, so we base our network topologies on currently deployed constellations, proposed future systems, and plausible extensions of existing systems. Table I gives some information about each scenario.
6.2.1 Earth/Moon/Mars
For the majority of this work, we focus on a representative network with communication between Earth, the Moon, and Mars. We have designed this network to match existing proposals for future interplanetary networks, such as NASA’s “LunaNet” proposal, in which relays enable communication between nodes on/orbiting Earth and the Moon [israelLunaNet2020].
[…]
We evaluate this network topology with 4 different starting configurations, changing the start time to ensure we cover every state of the relay links at the start of the experiment. In cases where the PKI system is not centralized, we also repeat the revocation experiments, changing the network segment from which the revocation originates.
6.2.2 Earth Constellations
[…]
6.2.3 Federation of CubeSats
[…]
6.2.4 Federation of Constellations
Finally, we consider networks composed of existing constellations talking to one another. This allows us to form a multi-layer satellite network, with satellites in Low Earth Orbit (LEO), Medium Earth Orbit (MEO), and/or Geosynchronous Equatorial Orbit (GEO). This has been proposed as a solution to increase coverage, reliability, and performance in satellite networks [akyildizMLSR2002]. This architecture is rapidly approaching in single-operator settings, with networks like Starlink planning to add additional layers of satellites to provide connectivity at higher latitudes [spacex2020fcc]. Although MEO and GEO constellations tend to provide good global coverage by themselves, connecting via LEO satellites can reduce the power requirements for ground terminals – indeed, LEO has been recently explored as an option to expand cellular and 5G connectivity [darwishLEO2022, diUltraDense2019, starlinkStarlink2023].
We consider the following case studies:
  • LEO/LEO: Two constellations in LEO communicating with one another to provide shorter traffic paths and better coverage. In this paper we consider the Iridium and Starlink constellations talking to one another.
[…]
  • LEO/GEO: GEO satellites provide even wider coverage than MEO, at the cost of much higher latency. We consider Iridium connecting to the Viasat-3 GEO satellites, consisting of 3 satellites spaced out to provide global coverage.
  • LEO/MEO/GEO: Iridium, mPOWER, and Viasat-3 all working together.
In each of these cases the satellites can connect to the ground directly, with additional inter-layer links to provide connectivity between each layer.
6.3 Routing
Routing in DTNs and satellite networks is a difficult problem, and has been explored extensively in recent literature [grasicEvolution2011, leguayDTN2005, aranitiContact2015, sobinSurvey2016, xiaogangSurvey2016, alagozExploring2007]. Its particular difficulties stem from constantly changing network topologies and many-hop paths, requiring a routing algorithm that can react to changes more rapidly than update messages can traverse the network.
The topology of many interplanetary networks makes it easier to route traffic – the use of only a small number of relays between segments means traffic only needs to be routed within each segment. Messages bound for other planets simply need to route to the relay, which can handle routing for the remainder of the journey.
In this paper we do not employ any particular routing algorithm, considering this to be beyond the scope of this work. To prevent results from being affected by a particular choice of routing algorithm or strategy, we assume all nodes have access to optimal routing information at all times, including looking ahead to the future state of the network in store-and-forward routing. This provides similar results to source routing, in which a message’s route is computed at the source then simply forwarded along the computed path.
6.4 Traffic Modeling
Figure 2: State of the relays from Earth to the Moon and Mars over time. Gaps indicate that the link is not available.
As discussed in Section 5, we do not rely upon randomized traffic modeling, instead considering pairwise communication between every pair of non-relay nodes in the network. This provides us with a distribution for each performance metric, which can be further partitioned into network segments, enabling a comprehensive analysis of the network’s performance under each PKI configuration.
6.5 Simulator
To run these experiments, a network simulator is required – such a simulator must be able to handle arbitrary satellite constellations and interplanetary networks, scale to large numbers of nodes, and be easily extensible to support new protocols. We found in Section 3 that existing simulators were not suitable due to inefficiency, lack of support for the network configurations tested, and computational inefficiency. We therefore present the Deep Space Network Simulator (DSNS): a new event-based simulator for satellite networks with the required features for our experimental framework, providing a simple interface for actors in the simulation to add and respond to events in a queue, while simulating the mobility and connectivity of satellites. Unlike existing network simulators, DSNS can more easily scale to large numbers of nodes, due to its computational requirements governed only by the number of events generated by the simulation. It supports arbitrary network topologies, with mechanisms currently implemented for the simulation of Walker constellations or real satellites by using the SGP-4 orbital propagator, in addition to dynamic connectivity simulations based on line of sight and distance (or simply fixed links, depending on requirements).
DSNS is easily extensible to new protocols and layers of the network stack – we were thus only required to implement the components of the network stack necessary for these experiments. This enables us to run realistic simulations without the overhead of running the entire network stack, and makes it easy to switch between different protocols and configurations without requiring the entire simulation to be rewritten. DSNS is also written in a modular format, allowing components and protocols to be added and removed as desired. This enables new functionality to be added without requiring a deep understanding of the rest of the system. The modularity of DSNS also enables new functionality to be added without requiring a deep understanding of the rest of the system. On publication, the source code to DSNS will be made freely and openly available to aid future research.
We are only testing PKI protocols, so the full network stack is not implemented; to do so would raise further issues, as the satellite network ecosystem is sufficiently fragmented that such simulations would only cover a fraction of cases. We rely instead upon the message passing system in DSNS, building the behavior of each protocol as a set of rules for message creation and response. Messages are not acknowledged, resulting in behavior similar to UDP, SPP, or BP without acknowledgment. This gives the best-case results in each case, as acknowledging receipt of messages would add up to $1$ round-trip time to each exchange. We also do not consider retransmissions, as the protocols and mechanisms for discovering message loss and retransmitting are complex and varied, and could significantly change the results depending on implementation. Instead, messages are simply dropped if they fail to route to their destination, enabling us to see the underlying results without this additional complexity. To compare retransmission strategies, operators could add the retransmission time to the latency of dropped messages, or implement retransmission in DSNS directly.
Figure 3: Latency over time for an Earth/Moon/Mars interplanetary network passing messages between nodes in each segment. Latency is plotted on a log-scale.
7 Results
In this section we evaluate the simulations and experiments designed in the previous sections.
7.1 Connectivity
TABLE II: Cost of establishing a new connection under each constellation topology. Standard OCSP is used in each case, with a single centralized CA.
Dropped (%) Latency (s) Establishment Overhead (s) Hops
Constellation Min Max Mean Std Max Mean Std Min Max Mean Std
CubeSat $32.970282$ $130.008137$ $599.009212$ $162.710239$ $79.344462$ $591.998755$ $101.278389$ $83.889008$ $2.000000$ $79.000000$ $29.443129$ $14.722231$
Iridium $0.000$ $0.009542$ $1.200939$ $0.148321$ $0.094143$ $0.170168$ $0.100080$ $0.036130$ $2.000000$ $28.000000$ $15.932813$ $4.000474$
Starlink $24.163603$ $0.004351$ $229.077722$ $4.147903$ $22.193722$ $1.115087$ $0.098494$ $0.118140$ $2.000000$ $179.000000$ $54.978528$ $30.263371$
LEO/LEO $0.000$ $0.004351$ $1.199617$ $0.157189$ $0.196855$ $1.113498$ $0.090434$ $0.085447$ $2.000000$ $177.000000$ $40.383088$ $27.654900$
LEO/MEO $0.000$ $0.009542$ $1.200939$ $0.148211$ $0.093608$ $0.170168$ $0.100080$ $0.036130$ $2.000000$ $28.000000$ $15.885049$ $3.966313$
LEO/GEO $0.000$ $0.009542$ $1.200939$ $0.148321$ $0.094143$ $0.170168$ $0.100080$ $0.036130$ $2.000000$ $28.000000$ $15.932813$ $4.000474$
LEO/MEO/GEO $0.000$ $0.009542$ $1.200939$ $0.148211$ $0.093608$ $0.170168$ $0.100080$ $0.036130$ $2.000000$ $28.000000$ $15.885049$ $3.966313$
Earth/Moon/Mars $0.046083$ $0.005875$ 6 248.1364456248.1364456\,248.1364456 248.136445 $393.704201$ 1 203.4802131203.4802131\,203.4802131 203.480213 4 988.7425084988.7425084\,988.7425084 988.742508 $110.495389$ $524.621493$ $2.000000$ $54.000000$ $18.015237$ $6.758139$
Figure 4: Distribution of latencies when establishing a new connection under each constellation topology, using standard OCSP with a single centralized CA.
Before looking at properties of the PKI systems, we must first understand the network topologies themselves. We are particularly interested in how the partitions of the network change over time: connectivity between interplanetary segments is possible if and only if relays are available. This is visualized in Figure 2: when the relays are up, the network joins into a single segment (although there is still high latency across the relays), and as the relays periodically pass behind the planet and become unavailable, the network is split into multiple segments.
We can also observe the latency of the network as it changes over time. Latency data for the terrestrial constellations can be found in Appendix C. We can also observe latency statistics from the connection establishment results in Table II: Starlink offers good coverage with low latency due to its use of many satellites low orbits, but as a result, data must travel over significantly more hops to reach its destination. We also note that Starlink struggles to connect to ground stations in higher latitudes due to the satellites’ orbits, resulting in dropped messages and high latency for these ground stations. The CubeSat network also struggles; with only $121$ satellites and no consistent pattern to their movement, there is insufficient coverage to provide consistent low-latency routing across the globe. However, connectivity is still possible, and it is likely that latency could be brought down with the addition of more satellites, or some relay satellites in MEO or GEO. The Iridium constellation provides good connectivity: higher orbits and fewer satellites result in lower hop counts at the cost of slightly greater latency compared to Starlink.
The federated networks of LEO, MEO, and/or GEO perform very similarly to their standard counterparts, with almost identical latencies. This is because long-distance messages will preferentially travel over the LEO satellites, particularly when there is no fixed cost for each hop taken, and when bandwidth and onboard processing power are not a concern.
TABLE III: Full results for the connection establishment scenario on the terrestrial and federated constellations, for each PKI configuration tested.
Configuration Dropped (%) Latency (s) Establishment Overhead (s) Hops
Distributed PKI Variant Min Max Mean Std Max Mean Std Min Max Mean Std
OCSP $0.046083$ $0.005875$ 6 248.1364456248.1364456\,248.1364456 248.136445 $393.704201$ 1 203.4802131203.4802131\,203.4802131 203.480213 4 988.7425084988.7425084\,988.7425084 988.742508 $110.495389$ $524.621493$ $2$ $54$ $18.015237$ $6.758139$
OCSP Stapling $0.000$ $0.005875$ 6 250.0082006250.0082006\,250.0082006 250.008200 $396.124230$ 1 207.4545201207.4545201\,207.4545201 207.454520 4 988.7500154988.7500154\,988.7500154 988.750015 $202.821676$ $845.982258$ $2$ $54$ $18.032991$ $6.663003$
OCSP Validator $0.000$ $0.005875$ 4 988.7500154988.7500154\,988.7500154 988.750015 $292.995173$ $870.657070$ 1 296.3290541296.3290541\,296.3290541 296.329054 $3.985649$ $70.135116$ $2$ $48$ $17.899088$ $5.451550$
CRL $0.000$ $0.005875$ 4 988.7500154988.7500154\,988.7500154 988.750015 $395.936068$ 1 152.2302341152.2302341\,152.2302341 152.230234 4 988.7425084988.7425084\,988.7425084 988.742508 $111.027417$ $383.545301$ $2$ $53$ $18.014782$ $6.602655$
CRL Broadcast $0.000$ $0.005258$ 3 985.0106603985.0106603\,985.0106603 985.010660 $284.908650$ $855.930365$ $2$ $24$ $6.046925$ $2.999405$
OCSP $0.000$ $0.005875$ 3 985.0626063985.0626063\,985.0626063 985.062606 $285.072152$ $855.964677$ $30.006808$ $0.163501$ $0.850528$ $2$ $37$ $16.485974$ $4.493723$
OCSP Stapling $0.000$ $0.005875$ 3 985.0106603985.0106603\,985.0106603 985.010660 $285.056255$ $855.893697$ $230.063601$ $0.370284$ $6.872744$ $2$ $36$ $16.559169$ $4.371214$
OCSP Validator $0.000$ $0.005875$ 3 985.0106603985.0106603\,985.0106603 985.010660 $284.965509$ $855.919599$ $10.068846$ $0.033743$ $0.096094$ $2$ $36$ $16.781173$ $4.243603$
OCSP Hybrid $0.000$ $0.005875$ 3 985.0106603985.0106603\,985.0106603 985.010660 $285.044413$ $855.885643$ $230.063601$ $0.153829$ $1.490624$ $2$ $41$ $16.051597$ $4.142714$
CRL $0.000$ $0.005875$ 6 248.0975096248.0975096\,248.0975096 248.097509 $303.028315$ $924.165239$ 6 248.0902466248.0902466\,248.0902466 248.090246 $18.119664$ $215.975843$ $2$ $51$ $16.722817$ $4.584613$
CRL Broadcast $0.000$ $0.005258$ 3 985.0106603985.0106603\,985.0106603 985.010660 $284.908650$ $855.930365$ $2$ $32$ $11.379067$ $3.296142$
(a) Establishment overhead of each PKI configuration.
(b) Overall hop count of messages.
Figure 5: Cost of initiating a connection in the Earth/Moon/Mars network, with a distributed CA in each network segment.
[…]
7.2 Connection Establishment
First we look at the cost of establishing a new connection under each configuration. To evaluate this, we ran $202$ simulations, aggregating the results into useful insights. [^7] The full statistics for each simulation can be found at deep-space-network-simulator.github.io.
Table II shows the latency, establishment overhead, and number of hops required to initiate communication under each network topology when using the OCSP protocol under its default configuration. We can also see the distribution of latencies in Figure 4. From these we gain some interesting insights about the effectiveness of various constellations. We can see that the CubeSat network has high latency, and struggles to route all messages – this is likely due to the satellites’ limited and patchy coverage. If simulations run for long enough it will likely eventually route messages, so such a network may be suitable for certain use cases, but making calls back to a centralized CA will incur significant latency costs. We can also see that Starlink achieves very low latencies on many messages but struggles to route some in a timely manner, due to the Phase 1 satellites struggling to reach ground stations at higher latitudes. This is mitigated in the LEO/LEO network, which can make use of Iridium to expand coverage – this federated network performs better than either constellation acting by itself, reducing the overhead of connection establishment. However, the other federated networks do not make significant use of the higher layer satellites, so the results are almost identical to Iridium acting alone. We also see that the interplanetary network performs significantly worse, with a very high time overhead and a large number of messages dropped. This is due to each message requiring a handshake with an Earth-based CA, even those received by nodes on other planets – this can be mitigated by distributing CAs throughout the network.
Figure 6: Coverage over time for a revocation, relative to the time at which the revocation was issued. Revocation originates from an Earth-based CA, under the base OCSP configuration with distributed CAs.
Next, we compare the different PKI configurations, focusing on the Earth/Moon/Mars network. We can see these summarized in Table III. In the centralized configuration of each protocol, latency and establishment overhead are both so great as to be unusable, since every connection establishment requires communication with an Earth-based CA. However, the distributed case shows a significant improvement in establishment overhead, with results comparable to those of the Earth-based constellations (Table II)! This reduction in overhead is further illustrated in Figure 5, enabling efficient communication across the entire network.
Although all the OCSP variants in the distributed configuration result in good performance, minimal establishment overhead is achieved by using the “Validator” configuration, in which the validity is stapled to the message in-transit. This is unsurprising, as it removes the need for a round trip during establishment. However, it does risk link congestion by sending a high volume of traffic via the CAs – to mitigate this, the “Hybrid” approach uses in-transit validation for messages traversing the relay links, and normal OCSP Stapling elsewhere. This retains the benefits of low overhead for cross-segment messages, providing better performance than OCSP Stapling while minimizing the risk of congestion.
Finally, we discuss the differences between CRL-based systems and OCSP. Using CRLs can reduce the number of messages sent – although the statistics may appear similar to OCSP, it is important to remember that the establishment exchange only needs to be done for the first message sent within the CRL’s expiry window, rather than being required for every new node initiating communication. Any subsequent messages do not require a CRL to be requested, as it is already cached. This can significantly lower the number of messages sent, at the cost of security – if a CRL is cached, then messages using a key revoked since the last CRL was sent will be accepted, until the next CRL is sent. Additionally, although the number of messages is lower, each CRL will be much larger than a single OCSP message as it contains the full list of revoked certificates. This may be mitigated through the use of techniques like “delta CRLs” to reduce the size of each message [cooperMore2000], although these risk opening up further threats through attackers blocking specific portions of the CRL from being delivered.
7.3 Key Revocation
For this scenario, we focus on the interplanetary Earth/Moon/Mars network, with Iridium as a reference terrestrial satellite network. We ran 1 04410441\,0441 044 simulations with different configurations, discussing the most interesting results in this paper. [^8] The full results are provided at deep-space-network-simulator.github.io.
Figure 6 illustrates the coverage time for a revocation under OCSP. We can see that the coverage time is often negative for segments in which the attacker is not present – this is due to the long distances between segments, enabling the revocation to reach the CA long before the attacker’s message traverses the relay link. As a result, only messages sent significantly before of the revocation will be accepted in that segment, resulting in a negative coverage time. This is explained further in Appendix D.
In contrast, when the attacker and revoker are located in the same segment, coverage time is primarily influenced by each node’s proximity to the CA, leading to rapid overall coverage with some nodes protected later than others. Our objective is to optimize the PKI configuration to achieve the earliest possible coverage in each scenario, minimizing damage within the speed-of-light constraints inherent in long-distance communication.
TABLE IV: Coverage results (time taken for revocation to reach all nodes) for simulations on the Iridium constellation, grouped by attacker location and CA from which the revocation originates.
Configuration Attacker Segment Coverage Time (s)
PKI Variant Earth Origin Space Origin
CRL Ground $0.022355$ $0.004453$
CRL Space $0.013384$ $0.000897$
CRL Broadcast Ground $0.065107$ $0.088384$
CRL Broadcast Space $0.073578$ $0.091556$
OCSP Ground $0.022355$ $0.004453$
OCSP Space $0.013384$ $0.000897$
OCSP Stapling Ground $0.022329$ $0.004423$
OCSP Stapling Space $0.013340$ $-0.041314$
OCSP Validator Ground $0.022329$ $0.004423$
OCSP Validator Space $0.013340$ $-0.041314$
We can see aggregated coverage information for a wider range of configurations in Tables IV (Iridium) and V (Earth/Moon/Mars). In these tables, the coverage time is the time after which every node in the segment is protected by the revocation. We can once again see a significant difference between the cases where the attack originates from the same segment as the revocation, and when it originates from a different segment – in the latter case it takes significantly longer for the segments to be covered. This is because the revocation does not always have time to propagate to the segment in which the attacker is present, and must instead race the attacker’s messages to the victim/CA.
The overall coverage statistics are quite similar between configurations, but there are some differences – the base OCSP and CRL configurations perform better in some cases, and OCSP Hybrid provides improvements over the Stapling and Validator configurations.
We have already established that for interplanetary networks a single centralized CA is not feasible due to high latency. Therefore, we do not include the centralized configurations in this analysis, although they are still present in the full dataset. In many of the centralized cases all the attacker’s messages can be rejected, but the latency overhead for legitimate communication is too high to be useful.
Similarly, we focus on the best-case scenario for the attacker, in which nodes already have the attacker’s certificate status cached, so they will not immediately request up-to-date revocation information from the CA. The results for when caching is not used can be found in the full dataset.
TABLE V: Results (time taken for revocation to cover all nodes) for simulations on the Earth/Moon/Mars constellation with a distributed CA. Attacker is in the Earth segment, results are grouped by revocation origin and victim segment.
Configuration Coverage Time (s)
PKI Variant Revocation Earth Moon Mars
OCSP Earth $0.016165$ $3.947191$ $4.834587$
OCSP Moon 1 002.5503731002.5503731\,002.5503731 002.550373 $-997.516107$ $682.893751$
OCSP Mars 2 474.3314712474.3314712\,474.3314712 474.331471 1 839.6034681839.6034681\,839.6034681 839.603468 2 466.847351-2466.847351-2\,466.847351- 2 466.847351
OCSP Stapling Earth $0.013790$ $2.242174$ $0.275946$
OCSP Stapling Moon 1 002.5473951002.5473951\,002.5473951 002.547395 1 002.9060461002.9060461\,002.9060461 002.906046 1 002.5881771002.5881771\,002.5881771 002.588177
OCSP Stapling Mars 2 474.3323302474.3323302\,474.3323302 474.332330 2 474.4243732474.4243732\,474.4243732 474.424373 2 474.7788722474.7788722\,474.7788722 474.778872
OCSP Validator Earth $0.013790$ $2.242174$ $0.275946$
OCSP Validator Moon 1 002.5473951002.5473951\,002.5473951 002.547395 1 002.9060461002.9060461\,002.9060461 002.906046 1 002.5881771002.5881771\,002.5881771 002.588177
OCSP Validator Mars 2 474.3323302474.3323302\,474.3323302 474.332330 2 474.4243732474.4243732\,474.4243732 474.424373 2 474.7788722474.7788722\,474.7788722 474.778872
OCSP Hybrid Earth $0.013790$ $2.242165$ $0.272925$
OCSP Hybrid Moon 1 002.5473951002.5473951\,002.5473951 002.547395 $996.104784$ 1 001.5937411001.5937411\,001.5937411 001.593741
OCSP Hybrid Mars 2 474.3323302474.3323302\,474.3323302 474.332330 2 472.4704172472.4704172\,472.4704172 472.470417 1 322.1102521322.1102521\,322.1102521 322.110252
CRL Earth $0.016165$ $3.786518$ $3.918009$
CRL Moon 1 002.5503731002.5503731\,002.5503731 002.550373 $-998.761400$ $682.074583$
CRL Mars 2 474.3373022474.3373022\,474.3373022 474.337302 1 839.5434381839.5434381\,839.5434381 839.543438 2 470.585934-2470.585934-2\,470.585934- 2 470.585934
CRL Broadcast Earth $0.087114$ $3.787188$ $3.774671$
CRL Broadcast Moon 1 002.6412911002.6412911\,002.6412911 002.641291 $-350.363271$ 1 004.8559661004.8559661\,004.8559661 004.855966
CRL Broadcast Mars 2 474.4187192474.4187192\,474.4187192 474.418719 2 162.1357812162.1357812\,162.1357812 162.135781 1 845.9740371845.9740371\,845.9740371 845.974037
7.3.1 Subscribe
Having noted that a large number of messages get accepted in cases where the attack and revocation originate from different segments of the network, we evaluate some potential methods of mitigating this threat. One such method requires OCSP CAs to register subscriptions for nodes that request a certificate status, forwarding any updates or revocations directly to that node. This is not possible for certain configurations (e.g. OCSP Stapling, as the CA does not know the destination of its stapled certificate status) but is good practice in supported cases, as caches of compromised certificates can be cleared sooner, reducing the window of opportunity for the attacker.
We can see in Table VI that this reduces the number of vulnerable nodes for the “OCSP Validator” configuration, but does not have a significant effect on the other configurations. However, this is still a good practice as it enables the system to converge faster following a revocation, minimizing the duration for which a compromised key can be used while still benefiting from caching.
TABLE VI: Results for the key revocation on Earth/Moon/Mars, when CAs subscribe nodes to updates on requested certificates using OCSP.
Accepted (%)
Configuration No Subscribe Subscribe
OCSP Variant Revocation Earth Moon Mars Earth Moon Mars
Earth $4.785156$ $39.583333$ $22.916667$ $4.785156$ $39.583333$ $22.916667$
Moon $99.316406$ $0.000$ $50.000$ $99.316406$ $0.000$ $50.000$
Mars $100.000$ $77.083333$ $0.000$ $100.000$ $77.083333$ $0.000$
Validator Earth $0.292969$ $6.250000$ $4.166667$ $0.292969$ $6.250000$ $4.166667$
Validator Moon $100.000$ $91.666667$ $95.833333$ $99.902344$ $91.666667$ $81.250000$
Validator Mars $100.000$ $100.000$ $91.666667$ $100.000$ $83.333333$ $91.666667$
Hybrid Earth $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
Hybrid Moon $100.000$ $0.000$ $52.083333$ $100.000$ $0.000$ $52.083333$
Hybrid Mars $100.000$ $87.500000$ $0.000$ $100.000$ $87.500000$ $0.000$
TABLE VII: Results for key revocation when the relay nodes are used as a firewall in the “OCSP Hybrid” configuration, compared to the base results.
Coverage Time (s)
Origin No Firewall Firewall
Attack Revocation Earth Moon Mars Earth Moon Mars
Earth Earth $0.013790$ $2.242165$ $0.272925$ $0.003222$ $-0.031606$ $-0.031606$
Earth Moon 1 002.5473951002.5473951\,002.5473951 002.547395 $996.104784$ 1 001.5937411001.5937411\,001.5937411 001.593741 1 002.5471841002.5471841\,002.5471841 002.547184 $-998.795207$ $679.667775$
Earth Mars 2 474.3323302474.3323302\,474.3323302 474.332330 2 472.4704172472.4704172\,472.4704172 472.470417 1 322.1102521322.1102521\,322.1102521 322.110252 2 474.3371102474.3371102\,474.3371102 474.337110 1 839.4629841839.4629841\,839.4629841 839.462984 2 470.585961-2470.585961-2\,470.585961- 2 470.585961
Moon Earth $714.547810$ $907.435565$ $5.739158$ 1 001.523576-1001.523576-1\,001.523576- 1 001.523576 $806.606563$ $-999.608869$
Moon Moon $1.057950$ $-2.149271$ $0.796097$ $-2.686824$ $-2.149602$ $-324.280611$
Moon Mars 2 641.3623472641.3623472\,641.3623472 641.362347 2 789.4254812789.4254812\,789.4254812 789.425481 $637.015396$ 1 472.7799091472.7799091\,472.7799091 472.779909 2 687.3231492687.3231492\,687.3231492 687.323149 3 472.081046-3472.081046-3\,472.081046- 3 472.081046
Mars Earth 2 319.6779302319.6779302\,319.6779302 319.677930 2 319.1343172319.1343172\,319.1343172 319.134317 2 334.3938142334.3938142\,334.3938142 334.393814 2 470.618104-2470.618104-2\,470.618104- 2 470.618104 2 466.968745-2466.968745-2\,466.968745- 2 466.968745 2 450.4834682450.4834682\,450.4834682 450.483468
Mars Moon 3 010.0593993010.0593993\,010.0593993 010.059399 3 009.6636633009.6636633\,009.6636633 009.663663 3 024.8689363024.8689363\,024.8689363 024.868936 1 468.127397-1468.127397-1\,468.127397- 1 468.127397 3 469.415524-3469.415524-3\,469.415524- 3 469.415524 3 128.6367113128.6367113\,128.6367113 128.636711
Mars Mars $3.741121$ $2.552029$ $0.127344$ $-0.003290$ $-633.609775$ $0.130869$
7.3.2 Firewall
Finally, we propose a new technique in which the relay nodes are used as a firewall to filter out messages which use revoked certificates. Under this configuration, the relay filters out all messages passing through which are making use of revoked keys, stopping them from ever reaching their destination or marking them as invalid (depending on implementation). This does not affect the message’s path, since all interplanetary traffic travels via the relays anyway, and in a number of cases (OCSP with Validators, OCSP Hybrid) the certificates are also validated by these relays. The technique has the potential to reduce the race between the revocation and the attacker’s messages when messages are waiting to be forwarded at a relay, allowing all of these messages to be dropped by the firewall instead. It also allows relays to drop messages from an attacker even if a valid certificate status has already been stapled to the message, reducing the reach of the attacker across long-distance relays.
The results of this can be seen in Table VII. We can see that this technique has a huge impact on the coverage time of the revocations, particularly when the attack and revocation originate from different segments – the attacker’s segment is largely unaffected, but the other segments in the network can be protected substantially earlier. This brings us closer to the expected best-case performance of revocations under distributed PKI, bounding the attacker’s reach to the smallest segment possible and covering the other segments in the network as early as possible. This technique also reduces load on the interplanetary relays, preventing denial of service by attackers using compromised keys to exhaust the bandwidth of relay links.
8 Discussion
Related works hypothesize that there is no “one size fits all” approach to PKI in satellite networks [viswanathanArchitecture2016, templinDTN2016]. Our results further confirm this: distributed OCSP with firewalls on each of the relays can provide high performance with fast revocation coverage, but some higher security applications may require centralized control, even at the cost of significantly increased latency. Our results enable satellite operators to make informed decisions about PKI implementation based on their specific needs and constraints: DSNS can be used to test the specifics of a PKI system against a network topology before launch and deployment, ensuring the desired balance between security and performance can be achieved. Additionally, our testing framework can be used in the future (alongside DSNS or another network simulator) to test newly proposed key management systems against a wide range of network topologies, establishing its performance and security characteristics.
When it comes to security, satellite operators are likely to have overlapping and occasionally incompatible needs, adding complexity when systems are networked together. We therefore propose a mixed-protocol system, in which key management systems are designed to work alongside one another. Both OCSP and CRLs are already highly intercompatible, as they are based upon existing internet technologies, and we have shown in our “OCSP Hybrid” protocol that mixed approaches are also possible in an interplanetary system. This would present the opportunity for organizations to choose PKI requirements on a per-device basis, permitting a high degree of control without sacrificing compatibility.
We must also consider the complexity of implementation of these technologies. Some works evaluated in Section 3 introduce new protocols; although this enables more significant changes, these new protocols have not been tested as extensively. By using protocols that have already been standardized and are in wide use across the terrestrial internet, we minimize the risk of security threats in the PKI system itself. We also benefit from the fact that standards already exist, so standardization bodies such as CCSDS do not need to build new standards from scratch, instead building off what already exists. This could involve further additions to the “BPSec” proposed standard [birraneBundle2022], adding new block definitions and behaviors to enable the new functionality. This offers the opportunity to accelerate the timeline for standardization, which is critical when the space industry is moving so quickly – a lack of standards for interplanetary communication when new systems are deployed could result in incompatibility between systems, making the goal of an interplanetary internet more difficult to realize.
8.1 Future Work
Alongside standardization efforts, some open questions remain which could be addressed in future work. To test cryptosystems and other implementation-specific details, DSNS could be extended to support bandwidth and storage simulation and message loss. This will enable regulators to figure out the specifics of how PKI protocols should be implemented to provide the fast connection establishment and effective revocations demonstrated in this paper, while also remaining within the computational and storage limits of satellite hardware. It will also enable demonstration of the concrete benefits of federated constellations, making use of a small number of powerful satellites to provide connectivity to a large number of devices while using LEO satellites for short-distance connections with low latency.
Future research may also consider testing “optimistic” protocols – that is, protocols which initially assume the certificate to be valid, rolling back the state if this is later shown not to be the case – or protocols like QUIC, which enable communication with zero round-trip-time setup [iyengarQUIC2021]. However, this raises additional security concerns: firstly, the system must be designed in such a way that state can be efficiently rolled back without losing other data or leaving artifacts behind. Many satellites have tight resource constraints, and it is vital that adversaries are prevented from exploiting these resources without proper authorization by opening temporary communication channels with invalid credentials. If these protocols are to be used, there must still be a mechanism for establishing contracts such that resources are not wasted on unauthorized users. Similarly, there is scope for evaluation of the scope and limitations for retroactive revocations, considering which actions can be safely rolled back without affecting general operation.
9 Conclusion
In this paper we have designed a framework for testing key management systems in satellite networks – this enables protocol designers to ensure their system is efficient and secure, and allows operators to optimize configurations for their network and specific needs. We have tested this framework against terrestrial PKI techniques applied to interplanetary networks using a novel network simulator. In contrast to current consensus, we have shown that terrestrial PKI can be used effectively in current and future satellite networks, ensuring efficient, low-latency connection establishment by distributing certificate authorities across the network.
We have also proposed new techniques building upon these, using OCSP Hybrid to combine concepts from existing OCSP configurations, ensuring revocations take effect as early as possible. Furthermore, by introducing a firewall at the interplanetary relays, the reach of an attacker across interplanetary links can be almost completely eliminated, preventing excess damage from the use of compromised keys and reducing the load on relay links.
By better understanding the impact of PKI configurations on the overall security of the system, we can ensure future interplanetary networks achieve maximal performance without sacrificing security.
Acknowledgments
The authors would like to thank Antonis Atlasis (ESA) for his insightful feedback which helped shape the paper, and armasuisse Science + Technology for providing funding and support for the project. Joshua was supported by the Engineering and Physical Sciences Research Council (EPSRC). Sebastian was supported by the Royal Academy of Engineering and the Office of the Chief Science Adviser for National Security under the UK Intelligence Community Postdoctoral Research Fellowships programme.
[…]
Appendix A Experiment Configuration
Figure 7: Locations of the ground stations used for each network topology in the simulations.
In the simulations built in Section 6, the terrestrial portion of the network contains 256 ground stations. These are randomly placed on Earth’s land surface, with the positions remaining fixed between simulations. The positions of the ground stations are visualized in Figure 7; exact coordinates will be included in the experiment configuration code released on publication (see Appendix E).
Appendix B Protocols
In this appendix we give sequence diagrams for each of the PKI techniques tested in this paper. These diagrams are shown for the distributed case, in which there are multiple CAs and authority is delegated. Note that the diagrams are also for nominal operation only – they do not illustrate behavior in the case of invalid signatures, caching, or revocations. OCSP Hybrid is not shown here, since behavior is identical to OCSP Stapling for communication within a segment, and identical to OCSP with Validators for communication between segments.
Full protocol specifications are in the source code for DSNS.
OCSP
ABCABSignedMessageCertificateQueryMessageCertificateStatusMessage
OCSP Stapling
ABCAACertificateQueryMessageCertificateStatusMessage StapledMessage ( SignedMessage, CertificateStatusMessage )
OCSP with Validators
ACARelayB ValidatorQueryMessage ( SignedMessage ) StapledMessage ( SignedMessage, CertificateStatusMessage )
CRL
ABCAASignedMessageCRLRequestMessageCRLDirectMessage
CRL Broadcast
Figure 8: Example of the race between attacker and CA, when an attack and revocation originate in different network segments.
ACAACRLBroadcastMessageCRLBroadcastMessageCRLBroadcastMessage
Appendix C Latency Results
Figure 9: Latency of messages between ground stations on opposite sides of the globe for non-federated terrestrial networks.
In this appendix we include additional results on the point-to-point latency of terrestrial constellations. Figure 9 illustrates how the latency changes over time for the Iridium and Starlink constellations, and the federated constellation of CubeSats evaluated in the paper.
Appendix D Revocation Coverage
In this appendix we provide further explanation of the “coverage time” results given for the revocation experiments in Section 7.3.
Figure 8 illustrates the race between an attacker’s message and a revocation, when they originate from different segments. Due to the vast distances involved, each segment is covered rapidly by its respective message, but cross-segment messages experience a significant delay. Consequently, the attacker must send their message well in advance of the revocation to ensure it is accepted, as the revocation message will reach the victim CA much more quickly than the attacker’s message. This can result in a negative coverage time, where messages sent before the revocation are rejected in the segment from which the revocation originates, while messages sent after the revocation are still accepted in the attacker’s segment.
Appendix E Data Availability
In the course of writing this paper, we ran a large number of network simulations, each of which produced substantial logs. These will be made fully available on publication via institutional research archives. Metrics for each simulation following processing can be found at deep-space-network-simulator.github.io.
Additionally, the source code for DSNS will be made available on publication as a Git repository, and released under a permissive license to enable future research. This will include the code used to configure and run the experiments, and process the results.